Solve each equation by factoring.
2p2+128=-32p

Answer by lenny460(1073)

Solve each equation by factoring.
2p2+128=-32p
I will assume 2p2 is 2p^2
Add 32p to each side of the equation.
2p^2 + 128 + 32p = -32p + 32p
Combine like terms:
-32p + 32p = 0
Therefore:
2p^2 + 128 + 32p = 0
Rearrange the above equation.
2p^2 + 32p + 128 = 0
Divide each term by 2
p^2 + 16p + 64 = 0
Factor the trinomial:
(p + 8)^2
Therefore:
(p + 8)(p + 8) = 0
Answer:
p = -8








Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com